OpenStudy (anonymous):

It is given that x-2 is a factor of P(x)=2x^(3)+px^(2)+qx+8. When P(x) is divided by x+1, the remainder is 27. Find (a) the values of p and q . (b) the remaining factors of P(x).

OpenStudy (shubhamsrg):

p(1) =0 p(-1)=27 does that help ?

OpenStudy (anonymous):

I know this

OpenStudy (anonymous):

p-q=21 ---(1) and the next step....

OpenStudy (shubhamsrg):

also make use of p(1) = 0

OpenStudy (anonymous):

ohhh....i type the question wrongly, it should be x-2 is a factor...

OpenStudy (shubhamsrg):

nevermind, then it'll be p(2) =0

OpenStudy (anonymous):

wait, let me calculate it..

OpenStudy (anonymous):

2p+q=12---(2)

OpenStudy (shubhamsrg):

12 or -12 ?

OpenStudy (anonymous):

-12

OpenStudy (shubhamsrg):

so you have 2 eqns, 2 variables, solve for p and q

OpenStudy (anonymous):

p=3, q=-18

OpenStudy (shubhamsrg):

yep

OpenStudy (shubhamsrg):

(Y)

OpenStudy (anonymous):

thanks

OpenStudy (shubhamsrg):

glad to help.

OpenStudy (anonymous):

and part b ... ??

OpenStudy (shubhamsrg):

oh I forgot! :P So, since you know p and q, you know P(x) right?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

long division?

OpenStudy (shubhamsrg):

One root of this you already know is 2

OpenStudy (shubhamsrg):

Well yes, long division would do the job

OpenStudy (shubhamsrg):

There is an alternative also though, you must be knowing sum of roots = -b/a, sum in pairs = c/a , product = -d/a ??

OpenStudy (anonymous):

ohh...after the long division, do i need to factorize it?

OpenStudy (anonymous):

ohhh, i hate sum and product of root /.\

OpenStudy (shubhamsrg):

Its up to you which method you want to follow and yes, after long division, you'll have to factorize

OpenStudy (anonymous):

thanks

OpenStudy (shubhamsrg):

WADECUM !

OpenStudy (anonymous):

(x-2)(2x-1)(x+4) ?

OpenStudy (shubhamsrg):

yep..seems right!